Circular No. 7073

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

                URL http://cfa-www.harvard.edu/iau/cbat.html

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1998fa IN UGC 3513 W. D. Li, E. Halderson, M. Modjaz, T. Shefler, J. Y. King, M. Pa-

penkova, R. R. Treffers and A. V. Filippenko, University of California at Berkeley, report their discovery of an apparent supernova in UGC 3513 during the course of the Lick Observatory Supernova Search (cf. IAUC 6627) with the 0.8-m Katzman Automatic Imaging Telescope (KAIT). The supernova was discovered and confirmed on unfiltered observations taken on Dec. 25.4 (mag ∼ 18.2) and 26.2 UT (mag ∼ 18.0). The new object is located at α = 6^{h}42^{m}51^{s}.51, δ = +41^{o}25′18′′.9 (equinox 2000.0), which is about 4′′.0 west and 3′′.8 north of the nucleus of UGC 3513. Inspection of a KAIT image of the same field on Dec. 22.4 showed that the supernova was present with an unfiltered mag of 18.7. A KAIT image on Dec. 18.3 (limiting magnitude ∼ 19.0) showed nothing at the position of the supernova.

1997 SZ_{10} AND 1996 TR_{66}

These transneptunian objects (MPEC 1998-Y09, 1998-Y28) have almost certainly been identified as librating about the 1:2 mean-motion resonance with Neptune. Using a selection of orbital elements by B. G. Marsden, H. F. Levison, Southwest Research Institute, and R. Malhotra, Lunar and Planetary Institute, have established cases that are likely to be stable for 10^{9} years. For 1997 SZ_{10} these cases have current osculating values a = 48.29 AU, e = 0.365, i = 11^{o}.8; 48.39, 0.367, 11^{o}.8; and 48.49, 0.370, 11^{o}.8. For 1996 TR_{66} representative cases have current a = 48.18 AU, e = 0.399, i = 12^{o}.4; 48.29, 0.401, 12^{o}.4; and 48.36, 0.402, 12^{o}.4. Furthermore, the objects currently have arguments of perihelion ω = 344^{o} and 311^{o}, respectively, and it appears that this quantity also librates, for each object by a few tens of degrees about a mean value near 320^{o}.

COMET C/1995 O1 (HALE-BOPP) G. J. Garradd, Tamworth, N.S.W., reports that there was a remarkable

3-mag brightening of the nuclear region of this comet between his CCD observations on Dec. 11 and 21. A. Pearce, Nedlands, W.A., observing visually with a 0.20-m reflector on Dec. 18.78 UT, found m_{1} = 11.3 and a coma 1′.0 across but noted that at low power the nuclear condensation contributed 90–95 percent of the light; at 380× the condensation was still relatively stellar and the coma almost nonexistent. On Dec. 19.78 the very prominent condensation made the outer coma very faint and ill-defined. By Dec. 25.75 the condensation had enlarged and dispersed through the coma, as in outbursts of comet 29P at comparable distances (r = 6–7 AU).

1998 December 26  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1998 CBAT  Brian G. Marsden